ELECTRONIC SPECTROSCOPY OF COBALT-NEON

\underline{T. C. CHENG}, S. HASBROUCK, M.A. DUNCAN, Department of Chemistry, University of Georgia, Athens, GA 30602-2556.

Co^{+}Ne was generated via laser vaporization in a pulsed supersonic expansion source, mass selected, and analyzed by visible photodissociation spectroscopy. An electronic band system was observed with an origin beginning at 13503 cm^{−1}. A progression of peaks beginning from the origin until the convergence limit can be seen, corresponding to the vibrational bands in the excited state of Co^{+}Ne. The excited state constants (we=124 cm^{−1}) were determined and the electronic cycle leads to a ground state binding energy (D_{0}=948 cm^{−1}). The ground state binding energy can be compared to other rare gas binding energies, which is correlated to the polarizability of the rare gas.